OpenStudy (gfdgdgdgfhdf):

State the property that is demonstrated: (9)(8 · 7) = (9 · 8)(7).

RhondaSommer (rhondasommer):

Well, what property do you think it is?

RhondaSommer (rhondasommer):

oh and \[\Huge\color{cyan}{\mathbb{WELCOME\\TO\\OPENSTUDY}}\]

OpenStudy (gfdgdgdgfhdf):

A. Distributive B. Associative C. Commutative D. Additive

RhondaSommer (rhondasommer):

You believe it is A?

OpenStudy (gfdgdgdgfhdf):

idk which it is

RhondaSommer (rhondasommer):

Distributive\[6(2*4)=36~then~6*2+6*4=36\]Associative\[(6)(2*4)=48~then~(2)(6*4)=48\]Communitave \[3*2=6~then~3*2=6\]Additive\[3+2=5~then~2+3=5\] Based on this, what is the answer?

OpenStudy (gfdgdgdgfhdf):

associative

RhondaSommer (rhondasommer):

good job :D thats right

OpenStudy (gfdgdgdgfhdf):

ty

RhondaSommer (rhondasommer):

make sure you hit the "best answer" button :D